# A General Framework for Joint Multi-State Models

### v3

@@ 3.3.1 Model Variants
Two baseline hazard conventions are common:
$\lambda_{0}^{k^{\prime}\mid k}(t\mid T_{im_{i}(t)})=\begin{cases}\lambda_{0}^{k^{\prime}\mid k}(t-T_{im_{i}(t)})&\text{(clock reset)},\\ \lambda_{0}^{k^{\prime}\mid k}(t)&\text{(clock forward)}.\end{cases}$
The clock-reset form models risk as a function of time spent in the current state, whereas the clock-forward form measures risk with respect to global time since study entry.
@@ 3.3.2 Likelihood Formulation
Assumptions for Likelihood Factorization A. Latent-level independence A1. Random effects $(b_{i})_{i}$ are mutually independent across individuals. B. Conditional independence within individuals B1. Longitudinal observations $(Y_{ij})_{ij}$ are mutually independent given $b_{i},X_{i}$. B2. Trajectories $(\mathcal{T}_{i}^{*})_{i}$ are mutually independent given $b_{i},X_{i}$. B3. Conditional on $b_{i},X_{i}$, the longitudinal and event processes are mutually independent. C. Censoring and process assumptions C1. Censoring times $(C_{i})_{i}$ are mutually independent and noninformative, i.e. $C_{i}\mathrel{\perp\!\!\!\perp}(Y_{i},\mathcal{T}_{i}^{*})\mid b_{i},X_{i}$. C2. Conditionally on $b_{i},X_{i}$, event trajectories satisfy the semi-Markov property 1: $p(\mathcal{T}_{i}^{*}\mid b_{i},X_{i})=\prod_{\ell\geq 0}p\left((T_{i,\ell+1},S_{i,\ell+1})\mid T_{i\ell},S_{i\ell},b_{i},X_{i}\right)$.
These grouped assumptions mirror those used in classical joint modeling [rizopoulos2012joint] and multi-state survival analysis [putter2007competing]: independence across subjects; the conditional independence structure linking the longitudinal and event submodels through shared random effects; noninformative censoring and (semi-)Markovian dynamics.
Let $\theta=(\gamma,Q,R,\alpha,\beta)$ denote the vector of model parameters. For subject $i$, we observe the longitudinal measurements $Y_{i}=(Y_{i1},\dots,Y_{in_{i}})$ and the event trajectory $\mathcal{T}_{i}=\{(T_{i\ell},S_{i\ell})_{0\leq\ell\leq m_{i}}\}$. The joint likelihood then factorizes as
$p(Y_{i},\mathcal{T}_{i}\mid X_{i},C_{i},\theta)=\int p(Y_{i}\mid b_{i},X_{i},\theta)\,p(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i},\theta)\,p(b_{i}\mid\theta)\,db_{i},$
where $b_{i}$ are individual random effects. This formulation generalizes the joint likelihoods of wulfsohn1997joint and ferrer2016joint to arbitrary multi-state event structures. Each part can then be explicitely expressed using Assumptions 3.3.2, yielding an expression very similar to that obtained by rizopoulos2012joint.
@@
$p(b_{i}\mid\theta)=\frac{(2\pi)^{-q/2}}{\det(Q)^{1/2}}\exp\left(-\tfrac{1}{2}b_{i}^{T}Q^{-1}b_{i}\right),$
@@
$p(Y_{i}\mid b_{i},X_{i},\theta)=\prod_{j=1}^{n_{i}}\frac{(2\pi)^{-d/2}}{\det(R)^{1/2}}\exp\left(-\tfrac{1}{2}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)^{T}R^{-1}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)\right),$
@@
$p(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i},\theta)=$ $\prod_{l=0}^{m_{i}-1}p\left((T_{i(l+1)},S_{i(l+1)})\mid b_{i},X_{i},(T_{il},S_{il}),\theta\right)$ (2)
$\exp\left(-\sum_{s:\,(S_{im_{i}},s)\in E}\Lambda_{i}^{k^{\prime}\mid S_{im_{i}}}(C_{i}\mid b_{i},X_{i},T_{im_{i}},\theta)\right).$
The proof of the expression of the Semi-Markov likelihood (2) is provided in Appendix B.
@@ 3.4 Model Selection and Information Criteria
To compare competing model specifications, we rely on information criteria based on the marginal likelihood as defined in Section 4. The Akaike Information Criterion (AIC) [akaike2003new] and the Bayesian Information Criterion (BIC) [schwarz1978estimating] are standard tools. While the AIC relies on an asymptotically unbiased estimator of the $\log$-likelihood, the BIC aims to identify the true model with great probability as the sample size grows.
@@
$\mathrm{AIC}\coloneqq-2\log\mathcal{L}_{\text{marginal}}(\hat{\theta};X,Y,\mathcal{T},C)+2k,$
where $k$ is the number of parameters, and $\hat{\theta}$ is a maximizer of the marginal likelihood.
The Bayesian Information Criterion is similarly defined but imposes a stronger penalty for model complexity, growing in $\log n$ where $n$ is the number of observations. Particularly in the context of mixed effects models, and even more so for multi-state joint models, the number of observations $n$ may vary from one definition to another, either the total number of repeated measurements or the number of individuals [delattre2014note]. To alleviate this problem, we recall the derivation of the BIC from the Laplace approximation [lebarbier2006introduction]. For a prior distribution $\pi$ on a set of models $\mathcal{M}$, the Laplace approximation approaches up to constant terms which could also be accounted to improve the approximation, the posterior probability given observed data $x$ defined in Section 4 by
@@
In both cases, one should aim at minimizing the chosen criterion.
4 Statistical Inference
The estimation of model parameters $\theta=(\gamma,Q,R,\alpha,\beta)$ relies on two likelihood formulations: the complete-data likelihood $\mathcal{L}_{\text{complete}}$, which includes latent variables, and the marginal likelihood $\mathcal{L}_{\text{marginal}}$, obtained by integrating them out
$\mathcal{L}_{\text{marginal}}(\theta;X,Y,\mathcal{T},C)=\int\mathcal{L}_{\text{complete}}(\theta;X,Y,\mathcal{T},C,b)\,db,$
where
$\mathcal{L}_{\text{complete}}(\theta;X,Y,\mathcal{T},C,b)=\prod_{i=1}^{n}p(Y_{i}\mid b_{i},X_{i},\theta)\,p(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i},\theta)\,p(b_{i}\mid\theta).$
It is to be noted that this integral is typically intractable in closed form due to the nonlinear nature of the model, necessitating numerical approximation methods such as Monte Carlo integration or Laplace approximation.
Several optimization methods adapted from the nonlinear mixed-effects literature can be applied, including Stochastic EM [kuhn2004coupling], Laplace approximation [wolfinger1993laplace] and Gauss–Hermite quadrature. These strategies are implemented in software such as JMBayes [jmbayes22024, rizopoulos2020package].
Another approach requiring mild regularity assumptions on the $\log$ marginal likelihood, but without the need for the model to belong in the exponential family, is to consider a stochastic gradient ascent scheme [caillebotte2025estimation, baey2023efficient] using Fisher’s identity and following the Robbins-Monro procedure [robbins1951stochastic], alongside with MCMC-based approximations such as Metropolis-Hastings and Hamiltonian Monte Carlo.
$\nabla_{\theta}\log\mathcal{L}_{\text{marginal}}(\theta;x)=\mathbb{E}_{b\sim p(\cdot\mid x,\theta)}\left(\nabla_{\theta}\log\mathcal{L}_{\text{complete}}(\theta;x,b)\right).$
This expectation is approximated using Monte Carlo samples from the posterior $p(\cdot\mid x,\theta)$, avoiding the need to evaluate the intractable marginal likelihood $\mathcal{L}_{\text{marginal}}(\theta;x)$. This formulation naturally supports stochastic gradient ascent with MCMC-based posterior sampling and is amenable to minibatch parallelization across individuals, ensuring the convergence to a critical point. The update rule is as follows:
Algorithm 1 Stochastic gradient ascent model inference
1:data $x$; initial parameter $\theta^{(0)}$; step sizes $(\eta_{t})_{t\geq 0}$ with $\sum_{t}\eta_{t}=+\infty$, $\sum_{t}\eta_{t}^{2}<+\infty$; batch size $K$; sampler $p_{t}^{\mathrm{MCMC}}$; stopping criterion.
2:
3:while not converged do
4:  Draw $b_{k}\sim p(\cdot\mid x,\theta_{n})\approx p_{n}^{\text{MCMC}}$.
5:  $\theta^{(t+1)}\leftarrow\theta^{(t)}+\frac{\eta_{t}}{K}\sum_{k=1}^{K}\nabla_{\theta}\log\mathcal{L}_{\text{complete}}(\theta^{(t)};x,b_{k})$.
6:end while
7:return $\hat{\theta}$
Preconditioning matrices may also be used, such as the Fisher Information Matrix in a natural gradient descent framework, or even specialized optimizers such as Adam [kingma2014adam], NAdam [dozat2016incorporating] or Adagrad [duchi2011adaptive]. In particular, the Adam optimizer will later be used in our numerical experiments.
@@ 5 Dynamic prediction
The simulation of individual trajectories using Algorithm 2 enables prediction in the multi-state joint modeling framework. Specifically, predictions are derived from simulated event trajectories, allowing for individualized risk assessment and forecasting of future states or event times based on a subject’s observed biomarker and event history. This approach generalizes classical dynamic prediction by leveraging the rich structure of multi-state processes and the joint distribution of longitudinal and event data.
Suppose we are interested in some functional of the true (unobserved) future trajectory,
@@
where $\mathbb{Y}$ is some generic space.
For example, one can think of the state taken by the individual $i$ at a time $u$. As is the case in traditional dynamic prediction however, we are only interested in predicting quantities or characteristics that depend on a yet unobserved future, given prior information, i.e. the trajectory up to some prediction time $t\leq u$ and observed longitudinal markers. Contrary to (single transition) joint models, the conditional probability distribution on $(\mathbb{R}\times\mathcal{S})^{\mathbb{N}}$ cannot be analytically derived.
Nonetheless, as seen in Algorithm 2, we are able to accurately simulate each trajectory. Therefore, a Monte Carlo estimation scheme may be devised under certain restrictions, where we estimate $\chi^{*}(\mathcal{T}_{i}^{*})$ by its expected value:
$\hat{\chi}_{i}=\coloneqq\mathbb{E}_{\mathcal{T}^{\prime}_{i}\sim\mathscr{L}(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{T}_{i},t)}\left(\chi^{*}(\mathcal{T}^{\prime}_{i})\right)\approx\frac{1}{B}\sum_{k=1}^{B}\chi^{*}(\mathcal{T}_{i}^{(k)}),\,\mathcal{T}_{i}^{(k)}\overset{\text{i.i.d.}}{\sim}\mathscr{L}(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{T}_{i},t).$
In general however, $\chi^{*}$ may very well depend on (countably) infinitely many transitions. As a result, since the proposed estiamtion relies on simulted samples, we require the quantity to depend only on a finite subset of transitions.
@@
Assume there exists $\chi:\bigcup_{n\geq 1}(\mathbb{R}\times\mathbb{S})^{n}\to\mathbb{Y}$ and $\tau_{i}$ a stopping time for the filtration $\mathcal{F}_{in}\coloneqq\sigma\left((T_{il},S_{il})_{0\leq l\leq n}\right)$ such that:
$\begin{cases}\tau_{i}<+\infty\text{ a.s.},\\ \chi^{*}(\mathcal{T}_{i}^{*})=\chi\left((T_{il},S_{il})_{1\leq l\leq\tau_{i}}\right)\end{cases}$
Essentially, given Assumption 1, with probability one we are guaranteed that the quantity of interested may be computed in a finite number of simulation steps. Indeed, the prediction algorithm for a single sample may be summarized as follows:
@@
4:while $n<\tau_{i}$ do
5:  Append simulated $(T_{in},S_{in})$ to $\mathcal{T}_{i}$ if $n>\operatorname{len}(\mathcal{T}_{i})$
6:  $n\leftarrow n+1$
@@
$\operatorname{depth}(G)\coloneqq\max_{(v_{0},v_{1},\dots,v_{k})}k,$
where $(v_{0},v_{1},\dots,v_{k})$ ranges over all directed paths in $G$.
@@
Let $G=(V,E)$ be a directed acyclic graph, $u\in\mathbb{R}$ be a fixed time, and $\mathbb{Y}=V$ be the set of possible states. Let $\tau_{i}=\inf\{n\in\mathbb{N}:T_{in}\geq u\text{ or }S_{in}\text{ is absorbing}\}$. Clearly, $\tau_{i}\leq\operatorname{depth}(G)<+\infty$. Then, take $\chi_{u}^{*}(\mathcal{T}_{i}^{*})=S_{i\tau_{i}}$, which corresponds to the state of the individual $i$ at time $u$. In particular, if $V=\{0,1\}$ and $E=\{(0,1)\}$, we recover the special case used for survival probability estimation in standard joint models [rizopoulos2011dynamic].

### v4

@@ 3.3.1 Model Variants
Two baseline hazard conventions are common and can be specified as
$\lambda_{0}^{k\to k^{\prime}}(t_{0},t)=\begin{cases}\lambda_{0}^{k\to k^{\prime}}(t_{0},t)&\text{(clock-reset)},\\ \lambda_{0}^{k\to k^{\prime}}(t)&\text{(clock-forward)}.\end{cases}$
The clock-reset form models risk as a function of time spent in the current state, whereas the clock-forward form measures risk with respect to global time since study entry. Common parametric baseline hazard specifications include the exponential ($\lambda_{0}(t)=\lambda$), the Weibull ($\lambda_{0}(t)=k\lambda^{k}t^{k-1}$), and the Gompertz ($\lambda_{0}(t)=ae^{bt}$) families, among others.
Furthermore, under the sojourn-time formulation, transition times are no longer constrained to be positive, since only the elapsed duration between successive transitions enters the baseline hazard function.
@@ 3.3.2 Likelihood Formulation
Assumptions for Likelihood Factorization A. Latent-level independence A1. Random effects $(b_{i})_{i}$ are mutually independent across individuals. B. Conditional independence within individuals B1. Longitudinal observations $(Y_{ij})_{ij}$ are mutually independent given $b_{i},X_{i}$. B2. Trajectories $(\mathcal{T}_{i}^{*})_{i}$ are mutually independent given $b_{i},X_{i}$. B3. Conditional on $b_{i},X_{i}$, the longitudinal and event processes are mutually independent. C. Censoring and process assumptions C1. Censoring times $(C_{i})_{i}$ are mutually independent and noninformative, i.e., $C_{i}\mathrel{\perp\!\!\!\perp}(Y_{i},\mathcal{T}_{i}^{*})\mid b_{i},X_{i}$. C2. Conditionally on $b_{i},X_{i}$, event trajectories satisfy the semi-Markov property 1 given $b_{i},X_{i}$.
These grouped assumptions mirror those used in classical joint modeling [rizopoulos2012joint] and multi-state survival analysis [putter2007competing]: independence across subjects, the conditional independence structure linking the longitudinal and event submodels through shared random effects, noninformative censoring, and semi-Markovian dynamics.
Let $\theta\coloneqq(\gamma,Q,R,\alpha,\beta)$ denote the vector of model parameters. For subject $i\in\llbracket n\rrbracket$, we observe the longitudinal measurements $Y_{i}=(Y_{i1},\dots,Y_{in_{i}})$ and the event trajectory $\mathcal{T}_{i}=\left\{(T_{i\ell},S_{i\ell})_{\ell\leq m_{i}}\right\}$. The joint likelihood then factorizes as
$p_{\theta}(Y_{i},\mathcal{T}_{i}\mid X_{i},C_{i})=\int p_{\theta}(Y_{i}\mid b_{i},X_{i})\,p_{\theta}(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i})\,p_{\theta}(b_{i})\,db_{i}.$
This formulation generalizes the joint likelihoods of wulfsohn1997joint and ferrer2016joint to arbitrary multi-state event structures. Each part can then be explicitly expressed using Assumptions 3.3.2, yielding an expression very similar to that obtained by rizopoulos2012joint.
@@
$p_{\theta}(b_{i})=\frac{(2\pi)^{-q/2}}{\det(Q)^{1/2}}\exp\left(-\tfrac{1}{2}b_{i}^{T}Q^{-1}b_{i}\right),$
@@
$p_{\theta}(Y_{i}\mid b_{i},X_{i})=\prod_{j=1}^{n_{i}}\frac{(2\pi)^{-d/2}}{\det(R)^{1/2}}\exp\left(-\tfrac{1}{2}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)^{T}R^{-1}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)\right),$
@@
$p_{\theta}(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i},\theta)=$ $\prod_{\ell=0}^{m_{i}-1}p_{\theta}\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell},S_{i\ell})\right)$ (3)
$\exp\left(-\sum_{j:(S_{im_{i}},j)\in E}\Lambda_{i,\theta}^{S_{im_{i}}\to j}(T_{im_{i}},C_{i}\mid b_{i},X_{i})\right).$
The proof of the expression of the Semi-Markov likelihood (3) is provided in Appendix B. Each term $p_{\theta}\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell},S_{i\ell})\right)$ can be computed using Equation 2.
@@
4 Statistical Inference and Model Selection
In this section, we outline a practical framework for statistical inference in the proposed multi-state semi-Markov joint model. We first derive an optimization scheme for estimating model parameters through a stochastic gradient ascent procedure, and then introduce model selection criteria to assess competing specifications. These methods will be illustrated and evaluated on simulated and real datasets in Section 6.
4.1 Stochastic Gradient Ascent
The estimation of model parameters $\theta=(\gamma,Q,R,\alpha,\beta)$ relies on two likelihood formulations: the complete-data likelihood $\mathcal{L}_{\mathrm{complete}}$, which includes latent variables, and the marginal likelihood $\mathcal{L}_{\mathrm{marginal}}$, obtained by integrating them out
$\mathcal{L}_{\mathrm{marginal}}(\theta;X,Y,\mathcal{T},C)=\int\mathcal{L}_{\mathrm{complete}}(\theta;X,Y,\mathcal{T},C,b)\,db,$
where $b\coloneqq(b_{1},\dots,b_{n})$, and owing to the independence across individuals
$\mathcal{L}_{\mathrm{complete}}(\theta;X,Y,\mathcal{T},C,b)=\prod_{i=1}^{n}p_{\theta}(Y_{i}\mid b_{i},X_{i})\,p_{\theta}(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i})\,p_{\theta}(b_{i}).$
It is to be noted that this integral is typically intractable in closed form due to the nonlinear nature of the model. Several optimization methods adapted from the nonlinear mixed-effects literature can be applied, including Stochastic EM [kuhn2004coupling], Laplace approximation [wolfinger1993laplace] and Gauss–Hermite quadrature, alongside MCMC-based approximations such as Metropolis–Hastings [hastings1970monte], Metropolis-within-Gibbs, Hamiltonian Monte Carlo [neal2011mcmc]. These strategies are implemented in software such as JMBayes [jmbayes22024, rizopoulos2020package].
Another approach requiring mild regularity assumptions on the $\log$ marginal likelihood, but without the need for the model to belong in the exponential family, is to consider a stochastic gradient ascent scheme [caillebotte2025estimation, baey2023efficient] using Fisher’s identity and following the Robbins-Monro procedure [robbins1951stochastic].
$\nabla_{\theta}\log\mathcal{L}_{\mathrm{marginal}}(\theta;x)=\mathbb{E}_{b\sim p_{\theta}(\cdot\mid x)}\left[\nabla_{\theta}\log\mathcal{L}_{\mathrm{complete}}(\theta;x,b)\right].$
This expectation is approximated using Monte Carlo samples from the posterior $p_{\theta}(\cdot\mid x)$, avoiding the need to evaluate the intractable marginal likelihood $\mathcal{L}_{\mathrm{marginal}}(\theta;x)$. To ensure the covariance matrices $Q$ and $R$ remain symmetric and positive-definite during optimization, we parametrize them using a $\log$-Cholesky decomposition of their inverse. By taking the logarithm of the diagonal elements of the Cholesky factors, we map the constrained covariance parameters to an unconstrained real-valued space, and avoid explicit matrix inversion for likelihood computation. This formulation naturally supports stochastic gradient ascent with Markov Chain Monte Carlo (MCMC)-based posterior sampling and enables parallelization across individuals, ensuring convergence to a critical point. The update rule is detailed in Algorithm 1.
Algorithm 1 Stochastic gradient ascent model inference
1:data $x$; initial parameters $\theta^{(0)}$; step sizes $(\eta_{t})_{t\geq 0}$ with $\sum_{t\geq 0}\eta_{t}=+\infty$, $\sum_{t\geq 0}\eta_{t}^{2}<+\infty$; number of parallel chains $K$; MCMC samplers $\{p_{k}^{\mathrm{MCMC}}\}_{k=1}^{K}$; stopping criterion.
2:
[…]
4:while not converged do
5:  For each $k\in\llbracket K\rrbracket$, draw $b_{k}\sim p_{\theta^{(t)}}(\cdot\mid x)\approx p_{k}^{\mathrm{MCMC}}$.
6:  $\theta^{(t+1)}\leftarrow\theta^{(t)}+\frac{\eta_{t}}{K}\sum_{k=1}^{K}\nabla_{\theta}\log\mathcal{L}_{\mathrm{complete}}(\theta^{(t)};x,b_{k})$
7:  $t\leftarrow t+1$
8:end while
9:return $\hat{\theta}$
Preconditioning matrices may also be used, such as the Fisher Information Matrix in a natural gradient ascent framework, or even specialized optimizers such as Adam [kingma2014adam], NAdam [dozat2016incorporating] or Adagrad [duchi2011adaptive]. In particular, the Adam optimizer will later be used in our numerical experiments.
@@ 4.2 Model Selection and Information Criteria
To compare competing model specifications, one typically relies on information criteria based on the marginal likelihood. In practice, competing models are compared by selecting the specification that yields the smallest value of the chosen information criterion. For instance, the Akaike Information Criterion (AIC) [akaike2003new] and the Bayesian Information Criterion (BIC) [schwarz1978estimating] are standard tools. While the AIC relies on an asymptotically unbiased estimator of the $\log$-likelihood, the BIC aims to identify the true model with high probability as the sample size grows.
@@
$\mathrm{AIC}\coloneqq-2\log\mathcal{L}_{\mathrm{marginal}}(\hat{\theta};X,Y,\mathcal{T},C)+2\dim\hat{\theta}$,
where $\dim\hat{\theta}$ is the number of parameters, and $\hat{\theta}$ is a maximizer of the marginal likelihood.
The Bayesian Information Criterion is similarly defined but imposes a stronger penalty for model complexity, growing in $\log n$ where $n$ is the number of observations. Particularly in the context of mixed effects models, and even more so for multi-state joint models, the number of observations $n$ may vary from one definition to another, either the total number of repeated measurements or the number of individuals [delattre2014note]. To alleviate this problem, we recall the derivation of the BIC from the Laplace approximation [lebarbier2006introduction]. For a prior distribution $\pi$ on a set of models $\mathcal{M}$, the Laplace approximation yields, up to constant terms (which could also be included to improve the approximation), the approximation of the posterior probability given observed data $x$ reads
@@
However, as previously noted, the marginal $\log$-likelihood
$\log\mathcal{L}_{\mathrm{marginal}}(\theta;x)=\log\int\mathcal{L}_{\mathrm{complete}}(\theta;x,b)\,db$
is generally intractable in mixed effects and joint models, since the integral with respect to the latent variables does not admit a closed form. Although the gradient of the marginal $\log$-likelihood can be efficiently estimated using Fisher’s identity, the marginal $\log$-likelihood itself cannot be evaluated directly.
One possibility is to estimate it using bridge sampling, which has been proven to provide consistent and accurate estimators of normalizing constants and marginal likelihoods based on MCMC output [meng1996simulating, gronau2017tutorial].
An alternative consists in approximating the posterior distribution $p_{\hat{\theta}}(b\mid x)$ by a Gaussian distribution, for instance via a Laplace approximation. Using the entropy identity,
$\log\mathcal{L}_{\mathrm{marginal}}(\hat{\theta};x)=\mathbb{E}_{b\sim p_{\hat{\theta}}(\cdot\mid x)}\left[\log\mathcal{L}(\hat{\theta};x,b)\right]-\mathbb{E}_{b\sim p_{\hat{\theta}}(\cdot\mid x)}\left[\log p_{\hat{\theta}}(b\mid x)\right],$
we can further leverage independence across individuals, so that the entropy term decomposes as
$\mathbb{E}_{b\sim p_{\hat{\theta}}(\cdot\mid x)}\left[\log p_{\hat{\theta}}(b\mid x)\right]=\sum_{i=1}^{n}\mathbb{E}_{b_{i}\sim p_{\hat{\theta}}(\cdot\mid x_{i})}\left[\log p_{\hat{\theta}}(b_{i}\mid x_{i})\right].$
Approximating each posterior $p_{\hat{\theta}}(\cdot\mid x_{i})$ by a Gaussian distribution with covariance matrix $\Sigma_{i}$, we obtain
$\log\mathcal{L}_{\mathrm{marginal}}(\hat{\theta};x)\approx\mathbb{E}_{b\sim p_{\hat{\theta}}(\cdot\mid x)}\left[\log\mathcal{L}(\hat{\theta};x,b)\right]+\sum_{i=1}^{n}\underbrace{\frac{1}{2}\left(\log\det\Sigma_{i}+q\left(\log(2\pi)+1\right)\right)}_{H_{i}},$
where $q$ denotes the dimension of the latent vector.
This approximation replaces the intractable posterior entropy with a sum of closed-form entropies of $q$-dimensional Gaussian distributions. Furthermore, each term $H_{i}$ depends only on its corresponding covariance matrix $\Sigma_{i}$, which can be easily estimated from MCMC samples of the latent variables using the empirical covariance estimator. The practical relevance of the proposed model selection criteria is discussed in Subsection 6.3.
@@ 5 Dynamic prediction
In this section, we turn to dynamic prediction within the multi-state framework. Although the prediction of longitudinal trajectories does not fundamentally differ from the standard nonlinear joint modeling setting, the key difference arises from the structure of the event process. Indeed, because the joint distribution of multi-state trajectories has no tractable closed-form expression, direct calculation of predictive quantities is generally infeasible.
To address this, we rely on simulation-based procedures which are rendered possible using Algorithm 3 (detailed in the Appendix). Specifically, predictions are derived from simulated event trajectories, allowing for individualized risk assessment and forecasting of future states or event times based on a subject’s observed biomarker and event history. This approach generalizes classical dynamic prediction by leveraging the rich structure of multi-state processes and the joint distribution of longitudinal and event data.
Let $i$ index a new individual. Suppose we are interested in some functional of the true (unobserved) future trajectory,
@@
where $\mathbb{Y}$ is some generic vector space.
For example, one can think of the state taken by the individual $i$ at a time $u$. As is the case in traditional dynamic prediction however, we are only interested in predicting quantities or characteristics that depend on a yet unobserved future, given prior information, i.e., the trajectory $\mathcal{S}_{i}(t)$ up to some prediction time $t\leq u$ and observed longitudinal markers $\mathcal{Y}_{i}(t)\coloneq\left\{(t_{ij},Y_{ij}):t_{ij}\leq t\right\}$. Contrary to (single transition) joint models, the conditional probability distribution on $(\mathbb{R}^{+}\times V)^{\mathbb{N}}$ cannot be analytically derived.
Nonetheless, as seen in Algorithm 3, we are able to accurately simulate each trajectory. Therefore, a single or double Monte Carlo estimation scheme may be devised under certain restrictions, where we estimate $\chi^{*}(\mathcal{T}_{i}^{*})$ by its conditional expected value (or other statistics depending on the posterior distribution, such as the median)
$\hat{\chi}_{i}^{(1)}\coloneqq\mathbb{E}_{\hat{\theta}}\left[\chi^{*}(\mathcal{T}^{\prime}_{i})\mid\mathcal{Y}_{i}(t),\mathcal{S}_{i}(t)\right]\approx\frac{1}{B}\sum_{k=1}^{B}\chi^{*}(\mathcal{T}_{i}^{(k)}),\,\mathcal{T}_{i}^{(k)}\overset{\mathrm{i.i.d.}}{\sim}\mathscr{L}_{\hat{\theta}}\left(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{S}_{i}(t)\right),$
where the samples $\{\mathcal{T}_{i}^{(k)}\}_{k=1}^{B}$ are generated using a MCMC algorithm targeting the distribution $\mathscr{L}_{\hat{\theta}}\left(b_{i}\mid\mathcal{Y}_{i}(t),\mathcal{S}_{i}(t)\right)$ in conjunction with Algorithm 3.
Another approach from Bayesian perspective treats the model parameters $\theta$ as random variables by imposing a prior distribution together with the training data $\mathcal{D}$ [rizopoulos2011dynamic], and thus in this case our new estimator may be defined
$\hat{\chi}_{i}^{(2)}\coloneqq\mathbb{E}_{\theta^{\prime}\sim\mathscr{L}(\theta\mid\mathcal{D})}\left[\mathbb{E}_{\theta^{\prime}}\left[\chi^{*}(\mathcal{T}^{\prime}_{i})\mid\mathcal{Y}(t),\mathcal{S}_{i}(t)\right]\right],$
and approximated by a double Monte Carlo scheme accordingly where $\mathscr{L}(\theta\mid\mathcal{D})$ is the posterior distribution of $\theta$.
In the case of longitudinal values, the procedure simplifies: for each draw of $b_{i}$, one usually predicts a single deterministic functional by substituting $\chi$ with $h(\cdot,\psi_{i})$, which already corresponds to the conditional mean of $Y_{i}$ given $b_{i}$. However, in general multi-state prediction, $\chi^{*}$ may very well depend on (countably) infinitely many transitions. As a result, since the proposed estimation relies on simulated samples, we require the quantity to depend only on a finite subset of transitions.
@@
There exists $\chi:\bigcup_{n\geq 1}(\mathbb{R}^{+}\times V)^{n}\to\mathbb{Y}$ and $\tau_{i}$ a stopping time for the filtration $\mathcal{F}_{in}\coloneqq\sigma\left((T_{il},S_{il})_{\ell\leq n}\right)$ such that
[…]
Essentially, given Assumption 1, with probability one we are guaranteed that the quantity of interest may be computed in a finite number of simulation steps. Indeed, the prediction algorithm for a single sample may be summarized as in Algorithm 2.
Algorithm 2 Prediction algorithm
@@
4:while $\ell\leq\tau_{i}$ do
5:  Append simulated $(T_{i\ell},S_{i\ell})$ to $\mathcal{T}_{i}$
6:  $\ell\leftarrow\ell+1$
@@
$\operatorname{depth}(G)\coloneqq\max_{(v_{0},\dots,v_{k})}k,$
where $(v_{0},\dots,v_{k})$ ranges over all directed paths in $G$.
@@
Let $G=(V,E)$ be a finite directed acyclic graph, $u\in\mathbb{R}^{+}$ be a fixed time, and $\mathbb{Y}=\Delta(V)$ be the simplex of distributions over states. Let $\tau_{i}=\inf\{n\in\mathbb{N}:T_{in}\geq u\text{ or }S_{in}\text{ is absorbing}\}$. Clearly, $\tau_{i}\leq\operatorname{depth}(G)<+\infty$. For each state $k\in V$, let $\mathds{1}_{k}\in\mathbb{Y}$ denote the vector that is $1$ at entry $k$ and 0 elsewhere (a one-hot encoding of $k$). Then, we set $\chi_{u}^{*}(\mathcal{T}_{i}^{*})=\mathds{1}_{S_{i\tau_{i}}}\in\mathbb{Y}$ to represent the state occupied by individual $i$ at time $u$. In particular, if $V=\{0,1\}$ and $E=\left\{(0,1)\right\}$, we recover the special case used for survival probability estimation in standard joint models [rizopoulos2011dynamic].